# Towards Robust Dialogue Breakdown Detection: Addressing Disruptors in Large Language Models with Self-Guided Reasoning

### v1

@@ 4.2 Advanced Prompting Strategies
$\Pi_{\mathrm{FS}}(\mathcal{H}_{i},s_{i};\,k)\;=\;\bigl{\{}\,(\mathcal{H}_{1}^{\prime},s_{1}^{\prime},b_{1}^{\prime}),\dots,(\mathcal{H}_{k}^{\prime},s_{k}^{\prime},b_{k}^{\prime})\bigr{\}}\;\cup\;(\mathcal{H}_{i},s_{i})$
Within the FS category, we differentiate:
@@
  • 2-Shot Hard (2S-Hard): We instead use two challenging samples where the correct outcome is subtle (annotator agreement 60–70%). One example might be a conversation with some confusion that eventually recovers (almost a breakdown, but not quite), and another might show a user hesitant or mildly frustrated (not a clear-cut hang-up, but dialogue quality is low). We hypothesize exposing the model to ambiguous scenarios can improve its performance under uncertainty, hopefully teaching it to output more calibrated confidence in similar situations.
  • 4-Shot (4S): We combine the Easy and Hard exemplars—two from 2S-Easy and two from 2S-Hard—thus exposing the model to a broader range of scenarios. While 4S can yield higher accuracy, prompt length grows, which may overwhelm smaller models.
    In datasets without direct human annotations such as BETOLD, we select short (15-20 turns) vs. long (21-30) dialogues as “easier” vs. “harder” exemplars, then mimic the above FS setups accordingly. In addition, since the annotators did not provide their reasoning, we used Llama-3.3 70B to generate the step-by-step reasoning field given each exemplar, the annotators’ probability distribution (treated as a confidence score) and the decision label.
4.3 Self-Guided Reasoning Techniques
@@
where $\Pi_{\mathrm{AR}}$ instructs the model to “recall relevant past dialogues” similar to $(\mathcal{H}_{i},s_{i})$. The final prediction is then based on both $\mathcal{H}_{i},s_{i}$ and the newly generated analogies $\mathcal{A}_{i}$. This self-guided technique removes the need for hand-crafted exemplars, potentially offering more tailored guidance to each test case.
4.4 Our Proposed Reasoning Technique
Curricular Learning with Analogical Reasoning (CL+AR). Curriculum learning involves training or prompting the model with an order of tasks that progresses from easy to hard, mimicking the way humans learn. The model builds a foundation to tackle more difficult ones by mastering simpler dialogue tasks. We propose applying this idea to LLM reasoning: the model is first steered to solve easy ”proxy” queries related to the target query, then gradually given harder versions of the problem. The easy queries and solutions serve as stepping stones, forming a curriculum for the model’s chain of thought. In a dialogue context, we first ask the LLM to generate $\mathcal{A}_{i}^{(1)}$ (an easy analogous example), then $\mathcal{A}_{i}^{(2)}$ (moderate difficulty), and finally $\mathcal{A}_{i}^{(m)}$ (close to the given input’s complexity). This gradually “walks” the model from simpler to harder analogies. We then request the final decision for $(\mathcal{H}_{i},s_{i})$. Let $\mathcal{A}_{i}^{(1)},\dots,\mathcal{A}_{i}^{(m)}$ be the generated set of analogies. The overall prompt is:
$\Pi_{\mathrm{CL+AR}}(\mathcal{H}_{i},s_{i})=\{\mathcal{A}_{i}^{(1)},\,\mathcal{A}_{i}^{(2)},\,\dots,\,\mathcal{A}_{i}^{(m)},\,(\mathcal{H}_{i},s_{i})\}$
@@ 4.5 Deployment Architecture
    $s_{i}\;=\;\mathcal{G}_{\mathrm{assistant}}\!\bigl{(}\mathcal{H}_{i},\;u_{i}\bigr{)}$
@@
    $s_{i}^{*}\;=\;\mathcal{G}_{\mathrm{superior}}\!\bigl{(}\mathcal{H}_{i},\;u_{i}\bigr{)}$
@@
This multi-tiered strategy significantly reduces computational demands and ensures the most resource-intensive models are only utilized in critical scenarios, thereby enhancing both efficiency and user experience in conversational AI deployments.
@@
5 Experiments
We evaluate our proposed approach using three prominent dialogue breakdown detection datasets: English and Japanese tracks from DBDC5 and the BETOLD task-oriented dialogue dataset.
5.1 Dialogue Breakdown Datasets
Dialogue Breakdown Detection Challenge 5 (DBDC5). DBDC5, originating from the WOCHAT IWSDS 2020 workshop, serves as a widely recognized benchmark for dialogue breakdown detection research [12]. The test set contains annotated dialogues in both English (1,950 dialogues) and Japanese (2,672 dialogues), involving interactions between humans and conversational agents. Each dialogue turn is annotated by 15–30 human annotators, classifying system utterances into three categories: Breakdown, Non-Breakdown, and Possible Breakdown (an intermediate category indicating uncertainty). DBDC5 (and its predecessors DBDC3/4) have been important for research in identifying when dialogues go off-track and why.
Breakdown Expectation for Task-Oriented Long Dialogues (BETOLD). Introduced in 2022, BETOLD identifies breakdowns specifically in task-oriented service dialogues, such as customer support calls [19]. This dataset comprises 13,524 real human–agent phone dialogues, annotated based on whether the interaction ended in a ”late user-initiated hang-up or forward” (LUHF). LUHF instances indicate user frustration, leading to either terminating the call prematurely or requesting human intervention. Approximately 33% of the dataset represents breakdown cases. BETOLD dialogues utilize abstract representations of intents and entities derived from NLU/NLG annotations instead of raw textual data to maintain user privacy. We follow official train-test splits provided by dataset creators, treating BETOLD’s automatic labels as ground truth for evaluation purposes.
5.2 Data Preprocessing and Experimental Setup
Following prior works [9, 34], we preprocess dialogue data from DBDC5 and BETOLD datasets by converting annotations into binary labels (Breakdown/Non-Breakdown). Ambiguous annotations (’Possible Breakdown’) are consolidated into the Breakdown class. To accommodate privacy constraints, the BETOLD dataset utilizes intent and entity abstractions rather than raw utterances, potentially limiting the efficacy of specific prompting techniques. We parse these structured dialogues and present them in a text form like “System: Intent: X — Entities: Y” for each turn, thereby preserving the sequence and dialogue flow in anonymized form. Basic cleaning (removal of extraneous symbols, ensuring consistent turn indexing) is applied for both datasets. No additional data augmentation is used beyond what is provided in the datasets.
@@ 5.2 Data Preprocessing and Experimental Setup
5.3 Fine-tuning Parameters
@@ 5.4 Evaluation Metrics
We use accuracy and F1 score as our primary evaluation metrics for breakdown detection. Accuracy measures the overall correctness of breakdown vs non-breakdown predictions. F1 score (the harmonic mean of precision and recall) is especially important because the breakdown class is under‑represented; it evaluates how well models detect breakdowns without neglecting the non‑breakdown class. We therefore report the F1 score for both classes along with overall accuracy, in line with prior work on DBDC5 [32, 34]. This mirrors the official challenge metrics and provides continuity with previous state-of-the-art results. Metrics are computed on held-out test sets (DBDC5 evaluation sets and reserved BETOLD splits). We parse the model’s JSON-formatted output to extract the decision label. In cases where model outputs deviate from the required JSON format, instead of discarding these responses, we employ Llama-3.3 70B as an LLM-based judge to verify output alignment with ground-truth labels [44].
@@
On the DBDC English and Japanese tracks, we investigated the model’s overconfidence using the MSE of the annotators’ probability distribution for a dialogue and the LLM’s verbalized confidence in its decision.
5.5 Inference and Cost Considerations
@@
For a comprehensive study, we examine a diverse range of LLMs, as listed in Table 2. The models are ordered beginning with proprietary models from organizations such as OpenAI (GPT-3.5 and GPT-4) [1], Anthropic (Claude-3.5 Haiku and Claude-3.5 Sonnet v2) [2]. In contrast, open-source models from Mistral AI (Mixtral 8x7B and Mixtral 8x22B) [20], DeepSeek (DeepSeek-R1) [21] and Meta (Llama-3.1 8B, Llama-3.3 70B and Llama-3.1 405B) [3] offer greater transparency to model details, increased flexibility for customization, and are often more cost-effective. For inference, we route queries through the OpenRouter API  [^2] https://openrouter.ai, which provides a unified interface to multiple LLM endpoints. This approach allowed us to evaluate different models under a consistent framework and logging. However, using these hosted models incurs costs per token. In practice, Claude 3.5 Sonnet v2 is the most expensive, roughly on the order of \$5/M input tokens and \$15/M output tokens. We carefully designed prompts to be concise (especially for zero-shot and few-shot setups) to control token usage. Techniques like AR and CL+AR produce longer interactions since the model generates examples as part of its answer, which we note as a trade-off. They may improve reasoning but use more tokens, impacting cost and latency. All models use a temperature of 0, which controls the randomness of language model output; a top-p value of 0.9, known as nucleus sampling, which further influences randomness; and a maximum token limit of 2048, reflecting the context-length constraints of smaller LLMs such as Mixtral 8x7B and certain Llama variants.
6 Results and Discussion
@@
This section presents detailed performance analyses across three datasets: BETOLD (task-oriented dialogues) and DBDC5 (English and Japanese tracks). We evaluate multiple prompting strategies, proprietary and open-source models, and our Dialogue Disruption Monitor. Table 2 summarizes accuracy, F1 scores for Breakdown (B) and Non-Breakdown (NB) classes, and calibration performance (MSE) where applicable.
@@
[…]
Previous SOTA BERT+SSMBA 73.9 78.2
S2T2 77.9 82.4 76.7 75.4
@@
Table 2: Experimental results of proprietary and open-source models on dialogue breakdown datasets. MSE represents verbalized confidence mean-squared error. Bold = best, underlined = second‐best.
6.1 Main Findings
State-of-the-Art Results. Our results confirm significant progress beyond earlier benchmarks. On the DBDC5 English track, multiple closed and open-source modern LLMs surpass the prior best (77.9% accuracy from S2T2). For instance, Claude-3.5 Sonnet v2 attains up to 85.5% accuracy (AR prompt), matching Llama-3.3 70B with CL+AR (85.5%). This tie represents the new top performance on DBDC5 English. On the DBDC5 Japanese track, the best outcome—89.0% accuracy—comes from Claude-3.5 Sonnet v2 with CL+AR, slightly above the leading open-source DeepSeek-R1 model at 87.0%. Thus, for Japanese, the Anthropic model retains a slight edge. Nonetheless, the top open-source systems now perform competitively, typically within 1–3 points of the best closed-source counterpart.
@@ 6.1 Main Findings
In contrast, GPT-3.5 Turbo underperforms substantially on BETOLD and has the greatest accuracy swings falling near 41%-64.5%. It tends to misclassify borderline “near-breakdown” utterances or produce imbalanced predictions. A plausible explanation is that GPT-3.5 Turbo is more sensitive to how examples are presented; certain prompt structures lead to skewed confidence or confusion in distinguishing near-breakdown from non-breakdown scenarios.
Open-Source Models. Larger open-source models (Llama-3.3 70B, Llama-3.1 405B, Mixtral 8x22B, DeepSeek-R1) match or exceed closed-source baselines on DBDC5 English (80%-85%) and demonstrate competitive performance on DBDC5 Japanese. However, on BETOLD, performance variability is pronounced, ranging from 68.3% (Llama-3.1 8B, 2S-Hard) to 75.8% (DeepSeek-R1, 2S-Easy), reflecting difficulties generalizing from natural dialogues to abstract intent representations. Llama-3.1 405B, despite its size, does not consistently surpass the 70B model. Its best English-track accuracy hovers near 79%–81%, indicating marginal variances to prompting techniques.
Dialogue Disruption Monitor. Despite having fewer parameters, our fine-tuned Llama-3.1 8B model achieves competitive performance. On the DBDC5 English dataset, it attains an accuracy of 81.5%, surpassing several larger models. For the DBDC5 Japanese dataset, the model achieves a moderate yet stable accuracy of 67.9%, with balanced F1 scores across classes. Even without fine-tuning specifically for the BETOLD dataset, the model demonstrates improved generalization for dialogue breakdown detection, increasing accuracy from 60.2% in a ZS-baseline to 67.2%. On DBDC5 English, the model’s performance is comparable to its significantly larger variants while demonstrating better calibration scores. However, performance on the Japanese dataset remains limited due to insufficient Japanese training data. Future research could benefit from utilizing models such as Llama-3.1 Swallow 8B, which has undergone additional training on 200 billion tokens derived from the extensive Japanese web corpus (Swallow Corpus Version 2), as well as Japanese and English Wikipedia [45].
6.2 Impact of Prompting Strategies

### v2

@@ IV-B Advanced Prompting Strategies
$\Pi_{\mathrm{FS}}(\mathcal{H}_{i},s_{i};\,k)\;=\;\bigl{\{}\,(\mathcal{H}_{1}^{\prime},s_{1}^{\prime},b_{1}^{\prime}),\dots,(\mathcal{H}_{k}^{\prime},s_{k}^{\prime},b_{k}^{\prime})\bigr{\}}\;\cup\;(\mathcal{H}_{i},s_{i}).$
Within FS, we explore:
@@
  • 2-Shot Hard (2S-Hard): We instead use two challenging samples where the correct outcome is subtle (annotator agreement 60–70%). One example might be a conversation with some confusion that eventually recovers (almost a breakdown, but not quite), and another might show a user hesitant or mildly frustrated (not a clear-cut hang-up, but dialogue quality is low). We hypothesize that exposing the model to ambiguous scenarios can improve its performance under uncertainty and lead to more calibrated confidence scores in similar situations.
  • 4-Shot (4S): We combine the Easy and Hard exemplars, two from 2S-Easy and two from 2S-Hard, thus exposing the model to a broader range of scenarios. While 4S can yield higher accuracy, prompt length grows, which increases token consumption and may approach context limits for some models, especially with longer dialogues.
    In datasets like BETOLD where per-utterance breakdown rationales from annotators are unavailable, we select short (15-20 turns) vs. long (21-30 turns) dialogues as “easier” vs. “harder” exemplars, then mimic the above FS setups accordingly. In addition, since the annotators did not provide their reasoning, we used Llama-3.3 70B to generate the step-by-step reasoning field given each exemplar, the annotators’ probability distribution (treated as a confidence score) and the decision label.
IV-C Self-Guided Reasoning Techniques
@@
where $\Pi_{\mathrm{AR}}$ instructs the model to “recall relevant past dialogues” similar to $(\mathcal{H}_{i},s_{i})$. The final prediction is then based on both $\mathcal{H}_{i},s_{i}$ and the newly generated analogies $\mathcal{A}_{i}$. This self-guided technique removes the need for hand-crafted exemplars, offering more tailored guidance for each test case.
IV-D Our Proposed Reasoning Technique
Curricular Learning with Analogical Reasoning (CL+AR). Curriculum learning involves training or prompting the model with an order of tasks that progresses from easy to hard, mimicking the way humans learn. The model builds a foundation to tackle more difficult ones by mastering simpler dialogue tasks. We propose applying this idea to LLM reasoning: the model is first steered to solve easy “proxy” queries related to the target query, then gradually given harder versions of the problem. The easy queries and solutions serve as stepping stones, forming a curriculum for the model’s chain of thought. In a dialogue context, we first ask the LLM to generate $\mathcal{A}_{i}^{(1)}$ (an easy analogous example), then $\mathcal{A}_{i}^{(2)}$ (moderate difficulty), and finally $\mathcal{A}_{i}^{(m)}$ (close to the given input’s complexity). This gradually “walks” the model from simpler to harder analogies. We then request the final decision for $(\mathcal{H}_{i},s_{i})$. Let $\mathcal{A}_{i}^{(1)},\dots,\mathcal{A}_{i}^{(m)}$ be the generated set of analogies. The overall prompt is:
$\Pi_{\mathrm{CL+AR}}(\mathcal{H}_{i},s_{i})=\{\mathcal{A}_{i}^{(1)},\,\mathcal{A}_{i}^{(2)},\,\dots,\,\mathcal{A}_{i}^{(m)},\,(\mathcal{H}_{i},s_{i})\}.$
@@ IV-E Deployment Architecture
    $s_{i}\;=\;\mathcal{G}_{\mathrm{assistant}}\!\bigl{(}\mathcal{H}_{i},\;u_{i}\bigr{)}.$
@@
    $s_{i}^{*}\;=\;\mathcal{G}_{\mathrm{superior}}\!\bigl{(}\mathcal{H}_{i},\;u_{i}\bigr{)}.$
@@
This multi-tiered strategy significantly reduces computational demands and ensures resource-intensive models are only utilized in critical scenarios, thereby enhancing efficiency and user experience in conversational AI deployments.
@@
V Experiments
We evaluate our proposed approach using three dialogue breakdown detection datasets: English and Japanese tracks from DBDC5 and the BETOLD task-oriented dialogue dataset.
V-A Dialogue Breakdown Datasets
Dialogue Breakdown Detection Challenge 5 (DBDC5). DBDC5, originating from the WOCHAT IWSDS 2020 workshop, serves as a widely recognized benchmark for dialogue breakdown detection research [12]. The evaluation set contains annotated dialogues in English (1,950 dialogues) and Japanese (2,672 dialogues), involving interactions between humans and conversational agents. Each dialogue turn is annotated by 15–30 human annotators. The system utterances are classified into three categories: Breakdown, Non-Breakdown, and Possible Breakdown (an intermediate category indicating uncertainty).
Breakdown Expectation for Task-Oriented Long Dialogues (BETOLD). Introduced in 2022, BETOLD identifies breakdowns specifically in task-oriented service dialogues, such as customer support calls [19]. This dataset comprises 13,524 real human–agent phone dialogues, annotated based on whether the interaction ended in a “late user-initiated hang-up or forward” (LUHF). LUHF instances indicate user frustration, leading to either terminating the call prematurely or requesting human intervention. Approximately 33% of the dataset represents breakdown cases. BETOLD dialogues utilize abstract representations of intents and entities derived from NLU/NLG annotations instead of raw textual data to maintain user privacy. We follow official train-test splits provided by dataset creators, treating BETOLD’s automatic labels as ground truth for evaluation purposes.
V-B Data Preprocessing and Experimental Setup
Following prior works [9, 34], we preprocess dialogue data from DBDC5 and BETOLD datasets by converting annotations into binary labels (Breakdown/Non-Breakdown). Ambiguous annotations (‘Possible Breakdown’) are consolidated into the Breakdown class. To accommodate privacy constraints, the BETOLD dataset utilizes intent and entity abstractions rather than raw utterances, potentially limiting the efficacy of specific prompting techniques. We parse these structured dialogues and present them in a text form like “System: Intent: X — Entities: Y” for each turn, thereby preserving the sequence and dialogue flow in anonymized form. Basic cleaning (removal of extraneous symbols, ensuring consistent turn indexing) is applied for both datasets. No additional data augmentation is used beyond what is provided in the datasets.
@@ V-B Data Preprocessing and Experimental Setup
V-C Fine-tuning Parameters
@@ V-D Evaluation Metrics
We use accuracy and F1 score as our primary evaluation metrics for breakdown detection. Accuracy measures the overall correctness of breakdown vs non-breakdown predictions. F1 score (the harmonic mean of precision and recall) is especially important because the breakdown class is under‑represented; it evaluates how well models detect breakdowns without neglecting the non‑breakdown class. We therefore report the F1 score for both classes along with overall accuracy, in line with prior work on DBDC5 [32, 34]. This mirrors the official challenge metrics and provides continuity with previous state-of-the-art results. Metrics are computed on held-out test sets (DBDC5 evaluation sets and reserved BETOLD splits). We parse the model’s JSON-formatted output to extract the decision label. In cases where model outputs deviate from the required JSON format, instead of discarding these responses, we employ Llama-3.3 70B as an LLM-based judge to interpret the model’s textual output and extract the intended classification decision when the JSON format is not strictly adhered to.
@@
On the DBDC English and Japanese tracks, we investigated the model’s overconfidence using the MSE between the annotators’ probability distribution for a dialogue utterance and the LLM’s verbalized confidence in its corresponding decision.
V-E Inference and Cost Considerations
@@
For a comprehensive study, we examine a diverse range of LLMs, as listed in Table II. The models are ordered beginning with proprietary models from organizations such as OpenAI (GPT-3.5 and GPT-4[1], Anthropic (Claude-3.5 Haiku and Claude-3.5 Sonnet v2[2]. In contrast, open-source models from Mistral AI (Mixtral 8x7B and Mixtral 8x22B[20], DeepSeek (DeepSeek-R1[21] and Meta (Llama-3.1 8B, Llama-3.3 70B and Llama-3.1 405B[3] offer greater transparency regarding model details, increased flexibility for customization, and are often more cost-effective. For inference, we route queries through the OpenRouter API  [^2] https://openrouter.ai, which provides a unified interface to multiple LLM endpoints. This approach allowed us to evaluate different models under a consistent framework and logging. However, using these hosted models incurs costs per token. In practice, Claude 3.5 Sonnet v2 is the most expensive, roughly on the order of \$5/M input tokens and \$15/M output tokens. We carefully designed prompts to be concise (especially for zero-shot and few-shot setups) to control token usage. Techniques like AR and CL+AR produce longer interactions since the model generates examples as part of its answer, which we note as a trade-off. They may improve reasoning but use more tokens, impacting cost and latency. All models use a temperature of 0. While larger models support longer contexts, a uniform maximum token limit of 2048 was applied across all models to ensure consistent comparison, particularly reflecting the constraints of some smaller models and API limitations for longer AR/CL+AR generations.
VI Results and Discussion
@@
This section presents detailed performance analyses across three datasets: BETOLD (task-oriented dialogues) and DBDC5 (English and Japanese tracks). We evaluate multiple prompting strategies, proprietary and open-source models, and our Dialogue Disruption Monitor. Table II summarizes accuracy, F1 scores for Breakdown (B) and Non-Breakdown (NB) classes, and calibration performance (MSE) where applicable.
TABLE II: Results of proprietary and open-source models on dialogue breakdown datasets. Bold = best, underlined = second‐best.
@@
Accuracy F1(B) F1(NB) Accuracy F1(B) F1(NB) MSE Accuracy F1(B) F1(NB) MSE
Prev. SOTA S2T2 77.9 82.4 76.7 75.4
@@
VI-A Main Findings
State-of-the-Art Results. Our results confirm significant progress beyond earlier benchmarks. On the DBDC5 English track, multiple closed and open-source modern LLMs surpass the prior best (77.9% accuracy from S2T2). For instance, Claude-3.5 Sonnet v2 attains up to 85.5% accuracy (AR prompt), matching Llama-3.3 70B with CL+AR (85.5%). This tie represents the new top performance on DBDC5 English. On the DBDC5 Japanese track, the best outcome of 89.0% accuracy comes from Claude-3.5 Sonnet v2 with CL+AR, slightly above the leading open-source DeepSeek-R1 model at 87.0%. Thus, for Japanese, the Anthropic model retains a slight edge. Nonetheless, the top open-source systems now perform competitively, typically within 1–3 points of the best closed-source counterpart.
@@ VI-A Main Findings
In contrast, GPT-3.5 Turbo underperforms substantially on BETOLD, with accuracy ranging from 41% to 64.5%. It tends to misclassify borderline “near-breakdown” utterances or produce imbalanced predictions. A plausible explanation is that GPT-3.5 Turbo is more sensitive to how examples are presented; certain prompt structures lead to skewed confidence or confusion in distinguishing near-breakdown from non-breakdown scenarios.
Open-Source Models. Larger open-source models (Llama-3.3 70B, Llama-3.1 405B, Mixtral 8x22B, DeepSeek-R1) match or exceed closed-source baselines on DBDC5 English (80%-85%) and demonstrate competitive performance on DBDC5 Japanese. However, on BETOLD, performance variability is pronounced, ranging from 68.3% (Llama-3.1 8B, 2S-Hard) to 75.8% (DeepSeek-R1, 2S-Easy), reflecting difficulties generalizing from natural dialogues to abstract intent representations. Llama-3.1 405B, despite its size, does not consistently surpass its 70B counterpart. Its best English-track accuracy hovers near 79%–81%, indicating marginal variances to prompting techniques.
Dialogue Disruption Monitor. Despite having fewer parameters, our fine-tuned Llama-3.1 8B model achieves competitive performance. On the DBDC5 English dataset, it attains an accuracy of 81.5%, surpassing several larger models. For the DBDC5 Japanese dataset, the model achieves a moderate yet stable accuracy of 67.9%, with balanced F1 scores across classes. Our DBDC5-fine-tuned monitor, when evaluated on the BETOLD dataset (on which it was not trained), achieved 67.2% accuracy. This represents a 7% absolute improvement over the base Llama-3.1 8B model’s zero-shot performance (60.2%) on BETOLD, indicating effective transfer of breakdown detection capabilities to a new dataset and format. On DBDC5 English, the model’s performance is comparable to that of significantly larger Llama variants while demonstrating better calibration scores. However, performance on the Japanese dataset remains limited due to insufficient Japanese training data. Future research could benefit from utilizing models such as Llama-3.1 Swallow 8B, which has undergone additional training on 200 billion tokens derived from the extensive Japanese web corpus (Swallow Corpus Version 2), as well as Japanese and English Wikipedia [44].
VI-B Impact of Prompting Strategies